OpenStudy (song_of_the_sole):

i really need help on this one question have no clue how to do it at all!!

OpenStudy (pxrson):

No solution, they are parallel.

OpenStudy (will.h):

There are 2 ways we can solve for system of equations. Do u know them?

OpenStudy (will.h):

Alright. We have substitution method. And elimination method

OpenStudy (will.h):

And that guy is right they are parallel but here's the proof

OpenStudy (will.h):

-2x -y = 1 -4x -2y = 1 I will use elimination method as it is easier in this case. How? We will need to eliminate one of the variables and then solve for the other So we will eliminate the x variables by multiplying the 1st equation by -2 -2(-2x -y) = 1 Simplify it 4x + 2y = 1 And now type in the other equation -4x-2y = 1

OpenStudy (will.h):

Notice that if we get to simplify that system of equations 4x + 2y = 1 -4x -2y = 1 ---------------------- 0x 0y = 2 See that the variables have been eliminated

OpenStudy (will.h):

You welcome. And sure post it up

OpenStudy (will.h):

Can u post it in a new question

OpenStudy (will.h):

Alright nvm

OpenStudy (will.h):

Are you gonna post it in new post?

OpenStudy (song_of_the_sole):

yeah

OpenStudy (will.h):

Oh okay. Tag me when u do